OpenStudy (anonymous):

5-4y+x+9y=? help anyone

OpenStudy (anonymous):

-4y+9y+5+x=0 5y+5+x=0

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can You help me out with another one

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

yes i'll try :)

OpenStudy (anonymous):

6(a-b)-5(2a+4b)

OpenStudy (anonymous):

6a-6b-10a-20b=0 -4a-26b=0

OpenStudy (anonymous):

got that ???

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

,